OpenStudy (anonymous):

Solve. x times five equals forty x equals eight x equals thirty-five x equals forty-five x equals two hundred

OpenStudy (puzzler7):

\[x=\frac{40}{5}\]

OpenStudy (anonymous):

right

OpenStudy (xapproachesinfinity):

@puzzler7 Show the steps pls

OpenStudy (anonymous):

what do you mean

OpenStudy (xapproachesinfinity):

we are not supposed to give answers but help get answer

OpenStudy (puzzler7):

\[5x=40\]\[\frac{5x}{5}=\frac{40}{5}\]\[x=\frac{40}{5}\]

OpenStudy (xapproachesinfinity):

thanks!

OpenStudy (puzzler7):

Welcome.

OpenStudy (xapproachesinfinity):

cowgirl, you need the simplify the fraction

OpenStudy (xapproachesinfinity):

what is 40/5

OpenStudy (anonymous):

oh thanks guys

OpenStudy (puzzler7):

@xapproachesinfinity What I did at first wasn't giving an answer, was it?

OpenStudy (anonymous):

oh guys one more

OpenStudy (xapproachesinfinity):

well it was a bit of an answer anyway! So you need to show what you did and how you achieved that! that is what we are supposed to do here

OpenStudy (xapproachesinfinity):

ok purr continue with her pls!

OpenStudy (anonymous):

Solve. five p equals twenty p equals twenty-five p equals twenty p equals five p equals four i think it is a but i am not sure

OpenStudy (puzzler7):

I will.

OpenStudy (puzzler7):

\[5p=20\]\[\frac{5p}{5}=\frac{20}{5}\]\[p=\frac{20}{5}\]

OpenStudy (anonymous):

ok thanks

OpenStudy (puzzler7):

Welcome.